# Spiking Meets Attention: Efficient Remote Sensing Image Super-Resolution with Attention Spiking Neural Networks

### v2

@@ Abstract
Spiking neural networks (SNNs) are emerging as a promising alternative to traditional artificial neural networks (ANNs), offering biological plausibility and energy efficiency. Despite these merits, SNNs are frequently hampered by limited capacity and insufficient representation power, yet remain underexplored in remote sensing super-resolution (SR) tasks. In this paper, we first observe that spiking signals exhibit drastic intensity variations across diverse textures, highlighting an active learning state of the neurons. This observation motivates us to apply SNNs for efficient SR of RSIs. Inspired by the success of attention mechanisms in representing salient information, we devise the spiking attention block (SAB), a concise yet effective component that optimizes membrane potentials through inferred attention weights, which, in turn, regulates spiking activity for superior feature representation. Our key contributions include: 1) we bridge the independent modulation between temporal and channel dimensions, facilitating joint feature correlation learning, and 2) we access the global self-similar patterns in large-scale remote sensing imagery to infer spatial attention weights, incorporating effective priors for realistic and faithful reconstruction. Building upon SAB, we proposed SpikeSR, which achieves state-of-the-art performance across various remote sensing benchmarks such as AID, DOTA, and DIOR, while maintaining high computational efficiency. The code of SpikeSR will be available upon paper acceptance.
Index Terms:
Image super-resolution, spiking neural network, attention mechanism, remote sensing.
@@
Figure 1: The visualization of pixel intensity and neuron voltage in images under various degradation factors reveals important insights. The pixel intensity map illustrates that the high-frequency components of the image tend to be smooth, indicating a reduction in sharp details during progressive downsampling. Neuron intensity maps, derived from a LIF model, show that high-frequency details persist with drastic fluctuations, suggesting that the neurons remain in an active state.
I Introduction
@@
High-resolution remote sensing images (RSIs) contain fine-grained object structures and textures, which are critical for accurate interpretation in downstream tasks [1, 2, 3]. However, limited by the intrinsic resolution of airborne sensors, RSI can merely capture partial spatial details, resulting in suboptimal scene representation and visual quality. Image super-resolution (SR) aims to alleviate this problem by reconstructing high-resolution (HR) images from low-resolution (LR) observations [4, 5]. Despite this, SR remains a challenging ill-posed issue, as a degraded input may correspond to multiple plausible outputs.
@@
Early efforts rely on hand-crafted priors to tame the ill-posedness, e.g., nonlocal mean [6, 7] and gradient profile [8], but they are often trapped in limited performance and scalability. Recent advances in artificial neural networks (ANNs), e.g., CNNs and Transformers, have witnessed remarkable progress in SR with large-capacity models [9, 10, 11, 12, 13]. However, they often come with a trade-off of increased computational overhead and growing storage costs, making them less efficient in practical scenarios, particularly when reconstructing large-scale RSIs.
@@
More recently, brain-inspired spiking neural networks (SNNs), as the third generation of neural networks, have emerged as a promising alternative for energy-efficient intelligence [14, 15, 16]. Different from ANNs that encode features as continuous values, SNN can emulate biological communication with discrete spiking signals and propagate them by neurons, thus enjoying lower power consumption. As depicted in Fig. 1, our experiments reveal a novel finding that spiking neurons maintain an active learning state across LR RSIs, even in severely damaged textures. Specifically, we observed that degraded RSIs exhibit smoothed pixel intensities and obscured sharp details, posing a significant challenge to characterize high-frequency representations. In contrast, spiking signals retain drastic responses and pronounced spike rates, highlighting that neurons remain in an active learning state. This naturally arises a question: Can SNNs leverage their inherent properties to handle image degradation for efficient yet high-quality RSI SR?
@@
In fact, to effectively grasp complex and diverse spatial details in RSIs, the network must possess adequate capacity and representation power. Unfortunately, there are two critical challenges when adapting SNNs for SR tasks. Firstly, spiking activity in SNNs inevitably causes pixel-wise information loss, which hampered the representation capacity of SNNs, especially when the network deepens. This stems from the discrete nature of binary spiking signals, leading to undesirable spiking degradation problems [17, 18]. Secondly, SNNs remain constrained by suboptimal membrane potential dynamics, restricting effective exploration of global context during spiking communications. This necessitates a customized strategy to optimize membrane potentials, but is barely explored before.
Figure 2: FLOPs and PSNR performance comparison. The circle sizes represent the number of parameters. Our SpikeSR outperforms SOTA efficient SR methods with high efficiency. PSNR results are averaged on the AID, DOTA, and DIOR datasets.
@@
To address these limitations, we propose SpikeSR, an SNN-based framework inspired by human visual attention mechanisms, which can actively represent image degradation and modulate synaptic weights to focus on salient regions, which, in turn, regulate the spiking activity for improved capacity and representation power. Specifically, SpikeSR employs a concise yet effective spiking attention block (SAB) to optimize feature emphasis through spiking response dynamics, which integrates three key innovations: 1) the combination of CNN and SNN layers to mitigate information loss induced by discrete spiking activity; 2) introducing hybrid dimension attention (HDA) to recalibrate spiking response across both temporal and channel dimensions, facilitating a joint feature correlations learning; 3) accessing global self-similarity patterns in RSIs to infer spatial attention weights, incorporating effective priors for realistic and faithful reconstruction. Compared to state-of-the-art (SOTA) ANN-based efficient SR models, our SpikeSR demonstrates lower model complexity and superior performance, as shown in Fig. 2.
@@
  • We pioneer an attention spiking neural network for efficient SR of remote sensing imagery, providing a new perspective on developing efficient models in large-scale Earth observation scenarios.
@@
  • We devise a concise yet effective spiking attention block, which mitigates the information loss and regulates membrane potentials of spiking activity for improved representation of SNNs.
@@
  • Extensive experimental results on various remote sensing datasets demonstrate that our SpikeSR achieves competitive SR performance against state-of-the-art ANNs-based methods.
Figure 3: Overall network architecture of SpikeSR. The LR input is replicated along the temporal dimension and then processed through a convolution to extract shallow features. The core module of SpikeSR is spiking attention groups (SAGs), which employ spiking attention blocks (SABs) to capture deep spiking representations. Each SAB contains three main components: (1) Spiking convolution block (SCB), (2) Hybrid dimension attention (HDA), and (3) Deformable similarity attention (DSA). The fusion block (FB) aggregates the spatial-temporal sequences, and pixelshuffle is used to reconstruct the SR output.
II Related Work
@@
Deep Networks for SR. Inspired by the pioneering SRCNN [19], CNN-based SR methods have achieved remarkable progress, dominating the field for years. They mainly elaborated on the network design to tame the ill-posedness, with notable advances in residual connections [20, 21] and attention mechanisms [22, 23]. For instance, to grasp valuable self-similarity priors, self-similar attention [24] and non-local sparse attention [25] have been developed. However, these methods often suffer from high computational complexity due to exhaustive non-local modeling, making them less efficient in large-scale RSIs.
@@
Recently, transformer-based SR models have demonstrated impressive performance, benefiting from their ability to model long-range dependencies. IPT [26] first introduces Transformers in SR field, but requires massive parameters and laborious pre-training processes. SwinIR [27] effectively reduces the model size by partitioning the image into smaller windows when applying multi-head attention mechanisms, while maintaining favorable performance. Despite these advancements, advanced SR models are often trapped by rising computational overhead and growing storage costs, posing significant concerns in real-world applications, particularly in remote sensing scenarios.
@@
Efficient SR Models. To reduce computational budget, CARN [28] utilizes grouped convolutions and a cascading mechanism to improve the residual architecture. IMDN [29] progressively distills useful information during feature extraction and applies network pruning to further decrease complexity. FMEN [30] optimizes residual modules to accelerate inference. In Transformer-based SR methods, SPIN [31] enhances long-range modeling by combining self-attention with pixel clustering, facilitating interactions between superpixels. HiT-SR [32] expands the self-attention receptive field by applying different window sizes of hierarchal layers. Despite these successes, there is still room to further boost SR performance. Moreover, the potential of energy-efficient SNNs for SR tasks remains largely unexplored.
@@
Spiking Neural Networks. Recent advances in neuromorphic computing have shown the great potential of SNNs in computational efficiency and power as CNNs. Currently, SNNs have been successfully applied to various tasks, such as image classification [14, 33], object detection and tracking [34, 15], optical flow estimation [35, 16], etc.
@@
A common solution to build SNNs is converting pre-trained ANN models [36]. Li et al. [37] proposed a layer-wise calibration to minimize activation mismatch during conversion. Ding et al. [38] replaced ReLU with the rate norm layer, enabling direct conversion from a trained ANN to an SNN. Stockl et al. [39] used time-varying multi-bit spikes to better approximate activation functions. However, conversion-based methods face accuracy gaps and high latency due to extensive time-step simulations, resulting in increased latency and energy consumption.
@@
An alternative involves using agent gradient functions for continuous relaxation of non-smooth spike activities, enabling direct training via backpropagation through time. Lee et al. [40] treated membrane potential as a signal to overcome discontinuities, enabling direct training from spikes. Wang et al. [41] introduced an iterative LIF model and proposed spatiotemporal backpropagation (STBP) based on approximate peak activity derivatives. Later, Zheng et al. [42] proposed temporal delay batch normalization (tdBN), which significantly enhanced the depth of SNNs. To bridge the performance gap between ANNs and SNNs, some methods borrowed insights from CNNs, applying residual learning [18, 43] and attention mechanisms [44, 45] to SNNs. Nonetheless, there has been limited exploration of pixel-level regression tasks, such as SR.
III Proposed Method
@@
The architecture of the proposed SpikeSR is illustrated in Fig. 3, which mainly consists of spiking attention groups (SAGs). Before SAGs, we utilize a $3\times 3$ convolution to extract high-dimensional features from the LR input. These features are then processed through $m$ stacked SAGs to explore deeper representations. Each SAG includes $n$ spiking attention blocks (SABs), a spiking convolution block (SCB), and a residual connection. In the SCB, leaky integrate-and-fire (LIF) neurons [46, 47] are used to convert the inputs into binary spike sequences (i.e., 0 or 1). As shown in Fig. 3, the output of the LIF neuron is 1 when the membrane potential exceeds the threshold, and 0 otherwise. To optimize the membrane potential, we introduce hybrid dimension attention (HDA), which refines the spiking activity using an efficient temporal-channel joint attention [48]. Furthermore, the proposed deformable similarity attention (DSA) is employed to introduce global context for accurate SR. After the terminal SAG, a fusion block (FB) is utilized to convert the spike sequence features into continuous values. Finally, SpikeSR generates super-resolved output from the fused features by applying pixel-shuffle [49] and a $3\times 3$ convolutional layer.
III-A Spiking Attention Block
@@
As evidenced in Fig. 1, regions degraded by different factors exhibit noticeable fluctuations when encoded by LIF, highlighting pronounced firing spike rates of neurons. This provides robust and latent informative spiking cues from LR images. Unlike ANNs that encode images into continuous decimal values, SNNs use discrete binary spike values for neuronal communication, and thus demonstrated undesirable information loss [50, 44], resulting in limited capacity to represent degraded LR images. To address this, the design philosophy of the SAB is focused on leveraging CNNs and attention mechanisms to regulate membrane potentials, facilitating high-quality feature representation for SR, which in turn affects the spiking activity.
@@
Different from previous works that focus solely on separate temporal and channel modulation [17, 44, 51], SAB enables joint temporal-channel adjustment of the spike response with HDA, effectively achieving interdependencies between the temporal and channel scopes.
III-B Deformable Similarity Attention
Non-local self-similarity has been recognized as an effective prior for SR tasks [52]. However, existing non-local attention mechanisms are computationally expensive due to exhaustive non-local operations, which impedes their efficiency in large-scale RSIs. In contrast, the proposed DSA efficiently grasps complex self-similar patterns in RSIs at the patch level to infer intricate spatial weights. Then, we utilize the cross-attention (CA) paradigm to enhance long-range communication, facilitating the fusion of useful context.
The details of DSA are shown in Fig. 4. Since the object scale exhibits explicit diversity in RSIs, the input feature $F$ is downsampled using bilinear interpolation, forming a multi-scale feature pyramid. For clarity, we demonstrate this process by dividing the initial features into 9 patches. The final DSA exploits a cascaded patch division strategy [53]. Specifically, each patch is first average pooled to capture its spatial characteristics, then reshaped and subjected to self-similarity computation, yielding a similarity matrix. The final self-similarity scores are fused via matrix multiplication to enhance the multi-scale representation. The best-matching patch $\mathbf{\bar{P}}_{i}$} with $\mathbf{P}_{i}$} can be obtained by:

### v3

@@ Abstract
Spiking neural networks (SNNs) are emerging as a promising alternative to traditional artificial neural networks (ANNs), offering biological plausibility and energy efficiency. Despite these merits, SNNs are frequently hampered by limited capacity and insufficient representation power, yet remain underexplored in remote sensing image (RSI) super-resolution (SR) tasks. In this paper, we first observe that spiking signals exhibit drastic intensity variations across diverse textures, highlighting an active learning state of the neurons. This observation motivates us to apply SNNs for efficient SR of RSIs. Inspired by the success of attention mechanisms in representing salient information, we devise the spiking attention block (SAB), a concise yet effective component that optimizes membrane potentials through inferred attention weights, which, in turn, regulates spiking activity for superior feature representation. Our key contributions include: 1) we bridge the independent modulation between temporal and channel dimensions, facilitating joint feature correlation learning, and 2) we access the global self-similar patterns in large-scale remote sensing scenarios to infer spatial attention weights, incorporating effective priors for realistic and faithful reconstruction. Building upon SAB, we proposed SpikeSR, which achieves state-of-the-art performance across various remote sensing benchmarks such as AID, DOTA, and DIOR, while maintaining high computational efficiency. Code of SpikeSR will be available at https://github.com/XY-boy/SpikeSR.
@@
Figure 1: (a) The visualization of pixel intensity and neuron voltage in images under various degradation factors reveals important insights. The pixel intensity map illustrates that the high-frequency components of the image tend to be smooth, indicating a reduction in sharp details during progressive downsampling. Neuron intensity maps, derived from a LIF model maass1997networks ; fang2023spikingjelly , show that high-frequency details persist with drastic fluctuations, suggesting that the neurons remain in an active state. (b) FLOPs and PSNR performance comparison. The circle sizes represent the number of parameters. Our SpikeSR outperforms SOTA efficient SR methods with high efficiency. PSNR results are averaged on the AID, DOTA, and DIOR datasets.
@@
High-resolution remote sensing images (RSIs) contain fine-grained object structures and textures, which are critical for accurate interpretation in downstream tasks razzak2023multi ; dong2024building ; cornebise2022open . However, limited by the intrinsic resolution of airborne sensors, RSI can merely capture partial spatial details, resulting in suboptimal scene representation and visual quality. Image super-resolution (SR) aims to alleviate this problem by reconstructing high-resolution (HR) images from low-resolution (LR) observations wang2020deep ; yao2020cross . Despite this, SR remains a challenging ill-posed issue, as a degraded input may correspond to multiple plausible outputs.
@@
Early efforts rely on hand-crafted priors to tame the ill-posedness, e.g., nonlocal mean zhang2012single ; dian2017hyperspectral and gradient profile sun2008image , but they are often trapped in limited performance and scalability. Recent advances in artificial neural networks (ANNs), e.g., CNNs and Transformers, have witnessed remarkable progress in SR with large-capacity models sun2023spatially ; tai2017image ; chen2023activating ; chen2023dual ; zhou2023memory . However, they often come with a trade-off of increased computational overhead and growing storage costs, making them less efficient in practical scenarios, particularly when reconstructing large-scale RSIs.
@@
More recently, brain-inspired spiking neural networks (SNNs), as the third generation of neural networks, have emerged as a promising alternative for energy-efficient intelligence lan2023efficient ; yang2019dashnet ; paredes2019unsupervised . Different from ANNs that encode features as continuous values, SNNs can emulate biological communication with discrete spiking signals and propagate them by neurons, thus enjoying lower power consumption. As depicted in Fig. 1(a), our experiments reveal a novel finding that spiking neurons maintain an active learning state across LR RSIs, even in severely damaged textures. Specifically, we observed that degraded RSIs exhibit smoothed pixel intensities and obscured sharp details, posing a significant challenge to characterize high-frequency representations. In contrast, spiking signals retain drastic responses and pronounced spike rates, highlighting that neurons remain in an active learning state. This naturally arises a question: Can SNNs leverage their inherent properties to handle image degradation for efficient yet high-quality RSI SR?
@@
In fact, to effectively grasp complex and diverse spatial details in RSIs, the network must possess adequate capacity and representation power. Unfortunately, there are two critical challenges when adapting SNNs for SR tasks. Firstly, spiking activity in SNNs inevitably causes pixel-wise information loss, which hampers the representation capacity of SNNs, especially when the network deepens. This stems from the discrete nature of binary spiking signals, leading to undesirable spiking degradation problems yao2021temporal ; fang2021deep . Secondly, SNNs remain constrained by suboptimal membrane potential dynamics, restricting effective exploration of global context during spiking communications. This necessitates a customized strategy to optimize membrane potentials, but is barely explored before.
@@
To address these limitations, we propose SpikeSR, an SNN-based framework inspired by human visual attention mechanisms, which can actively represent image degradation and modulate synaptic weights to focus on salient regions, which, in turn, regulate the spiking activity for improved capacity and representation power. Specifically, SpikeSR employs a concise yet effective spiking attention block (SAB) to optimize feature emphasis through spiking response dynamics, which integrates three key innovations: 1) the combination of CNN and SNN layers to mitigate information loss induced by discrete spiking activity; 2) introducing hybrid dimension attention (HDA) to recalibrate spiking response across both temporal and channel dimensions, facilitating a joint feature correlations learning; 3) accessing global self-similarity patterns in RSIs to infer spatial attention weights, incorporating effective priors for realistic and faithful reconstruction. Compared to state-of-the-art (SOTA) ANN-based efficient SR models, our SpikeSR demonstrates lower model complexity and superior performance, as shown in Fig. 1(b).
@@
  • We pioneer an attention spiking neural network for efficient SR of RSIs, providing a new perspective on developing efficient models in large-scale Earth observation scenarios.
@@
  • We devise a concise yet effective SAB, which mitigates the information loss and regulates membrane potentials of spiking activity for improved representation of SNNs.
@@
  • Extensive experimental results on various remote sensing datasets demonstrate that our SpikeSR achieves competitive SR performance against SOTA ANN-based methods.
1 Related Work
@@
Deep Networks for SR. Inspired by the pioneering SRCNN srcnn , CNN-based SR methods have achieved remarkable progress, dominating the field for years. They mainly elaborated on the network design to tame the ill-posedness, with notable advances in residual connections vdsr ; edsr and attention mechanisms rcan ; han . However, these methods often suffer from high computational complexity, e.g., exhaustive non-local modeling lei2021hybrid ; nlsa , making them less efficient in large-scale RSIs.
@@
Recently, transformer-based SR models have demonstrated impressive performance, benefiting from their ability to model long-range dependencies. IPT chen2021pre first introduces Transformers in SR field, but requires massive parameters and laborious pre-training processes. SwinIR swinir effectively reduces the model size by partitioning the image into smaller windows when applying multi-head attention mechanisms, while maintaining favorable performance. Against transformer, Mamba-based SR methods achieve comparable global model capacity with linear complexity guo2024mambair ; 10817590 ; guo2025mambairv2 . Despite these advancements, advanced SR models are often trapped by rising computational overhead and growing storage costs, posing significant concerns in real-world applications, particularly in remote sensing scenarios.
@@
Efficient SR Models. To reduce computational budget, CARN carn utilizes grouped convolutions and a cascading mechanism to improve the residual architecture. IMDN imdn progressively distills useful information during feature extraction and applies network pruning to further decrease complexity. FMEN fmen optimizes residual modules to accelerate inference. In Transformer-based SR methods, SPIN spin enhances long-range modeling by combining self-attention with pixel clustering, facilitating interactions between superpixels. HiT-SR hitsr expands the self-attention receptive field by applying different window sizes of hierarchal layers. Despite these successes, there is still room to further boost SR performance. Moreover, the potential of energy-efficient SNNs for SR tasks remains largely unexplored.
@@
Spiking Neural Networks. Recent advances in neuromorphic computing have shown the great potential of SNNs in computational efficiency and power as CNNs yao5 ; yao6 . Currently, SNNs have been successfully applied to various tasks, such as image classification lan2023efficient ; Shi_2024_CVPR , object detection and tracking kim2020spiking ; yang2019dashnet , optical flow estimation lee2020spike ; paredes2019unsupervised , etc.
@@
A common solution to build SNNs is converting pre-trained ANN models 7280696 ; yao4 . Li et al. li2024error proposed a layer-wise calibration to minimize activation mismatch during conversion. Ding et al. ijcai2021p321 replaced ReLU with the rate norm layer, enabling direct conversion from a trained ANN to an SNN. Stockl et al. stockl2021optimized used time-varying multi-bit spikes to better approximate activation functions. However, conversion-based methods face accuracy gaps and high latency due to extensive time-step simulations, resulting in increased latency and energy consumption.
@@
An alternative involves using agent gradient functions for continuous relaxation of non-smooth spike activities, enabling direct training via backpropagation through time. Lee et al. lee2016training treated membrane potential as a signal to overcome discontinuities, enabling direct training from spikes. Wang et al. wang2023masked introduced an iterative LIF model and proposed spatiotemporal backpropagation based on approximate peak activity derivatives. Later, Zheng et al. zheng2021going proposed temporal delay batch normalization, which significantly enhanced the depth of SNNs. To bridge the performance gap between ANNs and SNNs, some methods borrowed insights from CNNs, applying residual learning fang2021deep ; 10428029 and attention mechanisms yao2023attention ; qiu2024gated to SNNs. Nonetheless, there has been limited exploration of pixel-level regression tasks, such as SR.
2 Method
@@
The architecture of the proposed SpikeSR is illustrated in Fig. 2, which mainly consists of SAGs. Before SAGs, we utilize a $3\times 3$ convolution to extract high-dimensional features from the LR input. These features are then processed through $m$ stacked SAGs to explore deeper representations. Each SAG includes $n$ SABs, a SCB, and a residual connection. In the SCB, leaky integrate-and-fire (LIF) neurons maass1997networks ; fang2023spikingjelly are used to convert the inputs into binary spike sequences (i.e., 0 or 1). As shown in Fig. 2, the output of the LIF neuron is 1 when the membrane potential exceeds the threshold, and 0 otherwise. To optimize the membrane potential, we introduce HDA, which refines the spiking activity using an efficient temporal-channel joint attention tjca . Furthermore, the proposed DSA is employed to introduce global context for accurate SR. After the terminal SAG, an FB is utilized to convert the spike sequence features into continuous values. Finally, SpikeSR generates super-resolved output from the fused features by applying pixel-shuffle espcn and a $3\times 3$ convolutional layer.
Figure 2: Overall network architecture of SpikeSR. The LR input is replicated along the temporal dimension and then processed through a convolution to extract shallow features. The core module of SpikeSR is SAG, which employs SABs to capture deep spiking representations. Each SAB contains three main components: (1) SCB, (2) HDA, and (3) DSA. The fusion block (FB) aggregates the spatial-temporal sequences, and pixelshuffle is used to reconstruct the SR output.
2.1 Spiking Attention Block
@@
As evidenced in Fig. 1, regions degraded by different factors exhibit noticeable fluctuations when encoded by LIF, highlighting pronounced firing spike rates of neurons. This provides robust and latent informative spiking cues from LR images. Unlike ANNs that encode images into continuous decimal values, SNNs use discrete binary spike values for neuronal communication, and thus demonstrate undesirable information loss kundu2021spike ; yao2023attention , resulting in limited capacity to represent degraded LR images. To address this, the design philosophy of the SAB is focused on leveraging CNNs and attention mechanisms to regulate membrane potentials, facilitating high-quality feature representation for SR, which in turn affects the spiking activity.
@@
Different from previous works that focus solely on separate temporal and channel modulation yao2021temporal ; yao2023attention ; song2024learning , SAB adheres to temporal-channel joint attention tjca to realize joint adjustment of the spike response in HDA, effectively achieving interdependencies between the temporal and channel scopes. More details of HDA can be found in the Appendix.
2.2 Deformable Similarity Attention
Non-local self-similarity has been recognized as an effective prior for SR tasks 10675360 . However, existing non-local attention mechanisms are computationally expensive due to exhaustive non-local operations, which impedes their efficiency in large-scale RSIs. In contrast, the proposed DSA efficiently grasps complex self-similar patterns in RSIs at the patch level to infer intricate spatial weights. Then, we utilize the cross-attention (CA) paradigm to enhance long-range communication, facilitating the fusion of useful context.
The details of DSA are shown in Fig. 3. Considering that the object scale exhibits explicit diversity in RSIs, the input feature $\rm F$ is downsampled using bilinear interpolation, forming a multi-scale feature pyramid. For clarity, we demonstrate this process by dividing the initial features into 9 patches. Following the setting in liu2023coarse , the final DSA exploits a cascaded patch division strategy. Specifically, each patch is first average-pooled to capture its spatial characteristics, then reshaped and subjected to self-similarity computation, yielding a similarity matrix. The final self-similarity scores are fused via matrix multiplication to enhance the multi-scale representation. The best-matching patch $\mathbf{\bar{P}}_{i}$} with $\mathbf{P}_{i}$} can be obtained by: